# Reranking Social Media Feeds: A Practical Guide for Field Experiments

### v1

@@ Abstract.
Social media plays a central role in shaping public opinion and behavior, yet performing experiments on these platforms and, in particular, on feed algorithms is becoming increasingly challenging. This article offers practical recommendations to researchers developing and deploying field experiments focused on real-time re-ranking of social media feeds. This article is organized around two contributions. First, we overview an experimental method using web browser extensions that intercepts and re-ranks content in real-time, enabling naturalistic re-ranking field experiments. We then describe feed interventions and measurements that this paradigm enables on participants’ actual feeds, without requiring the involvement of social media platforms. Second, we offer concrete technical recommendations for intercepting and re-ranking social media feeds with minimal user-facing delay, and provide an open-source implementation  [^1] https://github.com/StanfordHCI/FeedMonitor. This document aims to summarize lessons learned, provide concrete implementation details, and foster the ecosystem of independent social media research.
Figure 1. Diagram of the typical feed reranking steps using a browser extension. First, the extension identifies the posts that need to be rescored (optional if all posts are included). The ranking model assigns a new score to the posts, and finally, the updated data is shown in the browser. The modified feed can be enriched with custom widgets, such as surveys, to assess the impact of the intervention.
@@ 1. Introduction
Social media plays a critical role in forming public opinion and influencing both individual behaviors and societal outcomes (Weeks et al., 2017). Many rely on social media to follow the news and stay informed (Center, 2023). Central to this media consumption behavior is the feed, which selects and ranks the social media posts that users are exposed to. It is the primary interface through which users engage with content, and it is the lens they use to make sense of the world. Most major platforms, including Facebook, X (Twitter), and Reddit, rely on their feed to give users a centralized view of relevant content. However, in doing so, these platforms must make algorithmic curation choices (Backstrom, 2016). In response, researchers have called for special attention to the impact of these feed algorithms, for example the kinds of content and behaviors that they uprank or downrank. These design choices are of substantial interest because they may profoundly shape our experiences and beliefs. Of particular concern are feed algorithms that aim to maximize user engagement and inadvertently amplify undesirable content, which may exacerbate polarization (Rathje et al., 2021; Kubin and von Sikorski, 2021), misinformation (Valenzuela et al., 2019), and antisocial behaviors (Saurwein and Spencer-Smith, 2021). These questions have remained largely unanswerable to the group of people without personal access to hundreds of millions of users—a group that probably includes you if you are reading this article.
@@
Investigating the causal impact of the feed ranking in an ecologically valid way is challenging. Despite efforts to open-source some of the key components of the platforms’ ranking algorithms (Twitter, 2023), their “black box” nature, reinforced by the lack of data (Tromble, 2021), complicates attempts to evaluate their causal effects on individual users and society as a whole. While the “logic” behind the algorithm may be open source, the algorithms’ outputs inherently depend on closed machine learning models that score aspects of each post and determine what is visible in the feed and in what order. Another challenge is the increasingly limited access to posts that would be considered to be shown in users’ feeds, known as the user’s inventory. Incomplete access to the key components of the feed—algorithms, models, and inventory—inherently limits the validity of any investigation.
@@
This article aims to bridge this gap by offering a practical guide for conducting independent field experiments to investigate the impact of feed ranking in social media. The guidelines focus on facilitating the design of experiments that do not require the direct cooperation of the social media platform, enabling audits of feed algorithms and development of interventions that aim to improve the design of online social spaces. We describe how to implement experiments on the web interfaces of social media platforms using browser extensions, a feature available on all major browsers, by catching and customizing the network calls that populate the feed. We focus on specific interventions that this method supports, including feed re-ranking and content editing, and describe the relative advantages and disadvantages of each. We do not aim to exhaustively list all possible ways to implement a field experiment, but we do provide concrete building blocks for designing custom experiments.
@@
Figure 1 summarizes the main steps of a typical field experiment designed to customize the ranking of a social media feed with a browser extension. This guide focuses on designing interventions that do not affect the users’ typical experience on the platform, aiming to limit latency and any interface behavior that can disrupt the user experience. To achieve this goal, the browser extension (1) captures the feed, (2) applies any experimental transformations, then (3) inserts any custom user interface elements. First, it reads the feed to identify the posts that need editing according to the objective of the experiment. For example, the experiment may focus on all the content in the feed, or it may target specific posts such as politics or misinformation. Then, the extension can apply the desired intervention to the selected posts (or those newly inserted), such as changing their rank position, rewriting their text, or editing the social metrics. Finally, the extension can change the design of the webpage or add widgets that enhance the interface and enable new user interactions. Such modifications enable researchers to design custom user interactions or assess the effect of a specific intervention by adding targeted surveys.
@@
This article is organized as follows. In Sec. 2, we summarise the state-of-the-art and highlight studies closely related to our work. In Sec. 3, we present opportunities and challenges in changing the feed ranking with a browser extension, followed by an overview of possible approaches to measuring the outcomes of the field experiment. Finally, Sec. 5 provides key implementation details to develop the feed experiment, ranging from ensuring low-effort user registration to intercepting and editing the feed. We conclude with a discussion of implications.
@@ 2. Background
In this section, we summarize previous efforts to study social media through experiments and position this article in the current literature. On-platform feed experiments are the gold standard for inferring the causal effects of interventions on social media, as they test interventions in the same setting where users would typically encounter them, maximizing ecological validity. An alternative to running experiments on the platform is creating experimental clones of the platform and simulating a social media environment (Voggenreiter et al., 2023; Epstein and Lin, 2022; Jia et al., 2023). Developing a custom social media feed simulator is a meaningful approach, as it allows greater control over the experimental setup, such as the posts visualized by the user and features offered by the websites. This full control of the feed has significant advantages because participants can be exposed to the same content, potentially leading to a more controlled experiment that may help measure small effects. However, the downside is that the individual participants’ online experience is generally hard to reproduce, and the generalizability of the findings may be questioned. The participants may be exposed to a different interface and a different type of posts they typically consume, they may miss realistic social signals and dynamics, or they may be required to engage in activities not aligned with their daily habits. The requirements to engage with a custom platform may also exacerbate phenomena such as the Hawthorn effect, potentially altering the participants’ behavior that does not fully generalize to a real-world scenario.
Previous work shows the effectiveness of field experiments on social media and that they can bring valuable findings. Some examples include investigating the effect of changing Facebook’s feed ranking from engagement-based to chronological during a US election (Guess et al., 2023), the impact and propagation of emotions (Kramer et al., 2014), the amplification of political discourse (Huszár et al., 2022), the effect of perspective-taking on political polarization (Saveski et al., 2022) or using ads to reduce misinformation (Lin et al., 2024) and support charitable donations (Jilke et al., 2019).
However, designing a field experiment with live intervention on social media feeds presents many technical challenges, especially when done without the involvement of the platforms (Bode et al., 2020; Mosleh et al., 2022; Guess, 2021). The present work aims to support researchers with practical recommendations for designing field experiments focused on feed ranking.
@@
While many field experiments of interest only require changing participants’ feeds, some experiments remain out of scope for this approach. For example, this approach cannot perform network bucket testing or other experimental protocols that require changing entire groups of users’ experiences together (Ugander et al., 2013; Saveski et al., 2017), e.g., to understand equilibrium effects rather than individual effects. Our approach also cannot easily reach the user’s full inventory unless the experimenter implements a separate application layer to acquire it directly from the platform; typically, in the context of this paper, we will be focused on changes to the items that are already chosen to be shown to the user.
@@ 2.1. Key concepts
Browser extensions. This guide focuses on modifying social media feeds using Browser extensions. Browser or web extensions are add-ons typically available via the browser-specific store, designed to enhance the browser’s functionality. They support developers in customizing the browser’s behavior and modifying websites’ content. This capability is advantageous in developing custom social media feed interventions, enabling design control trials or A/B tests. From the perspective of the experimental participants, the only technical prerequisite is to install the extension and accept the permissions. With authorization, browser extensions can modify the pages of interest to achieve the desired intervention. These modifications to the web pages range from updating the content and its visual appearance to the website’s behavior. Sec. 5 provides an overview of implementation details.
Platform Inventory. In this guide, we refer to the full database of candidate posts used by the feed algorithm as the participant’s inventory. This is a crucial concept because, depending on the platform of interest, the level of access to the candidate set can enable or limit some interventions. When the feed is organized chronologically, the algorithm can sort by time the posts of the relevant users (e.g., friends or followed users) and return the most recent ones. When the feed is algorithmically curated, the algorithm needs to evaluate a large set of posts by users inside and outside the focal user’s first-degree social network. This process typically relies on multiple steps. For instance, Twitter/X’s “For you” algorithm, as described in the open-sourced implementation (Twitter, 2023), first selects content using a high-recall retrieval system that scans a large amount of data which is then scored and ranked by a more refined model.
@@ 3. Feed Ranking
This section summarises the potential of browser extensions to customize social media feeds. The re-ranking can be a complete shuffle of the feed based on a new objective or more focused, for example up-ranking or down-ranking content the researchers believe should receive more or less exposure. In this guide, we focus on these two use cases, up-ranking and down-ranking, where the content of interest is respectively demoted or promoted in the feed, but these recommendations can be adapted for different types of rank adaptations. These edits can be achieved by combining up-ranking and down-ranking operations, which constitute the building blocks of a feed intervention.
A key functionality of the system handling the experiment is identifying relevant content and scoring it according to the metric of interest. Depending on the research goal, this capability can be achieved with different levels of complexity, ranging from simple keyword matching to more complex logic relying on external AI services like large language models (LLMs). For example, feed posts can be scored and rearranged to demote hostile political content or prioritize posts reflecting positive emotions.
These manipulations offer a unique opportunity to investigate the causal effects of variations of the rank objective in a realistic scenario. We summarise the challenges and limitations of both implementations. Additionally, a browser extension can edit the content of the posts. We conclude this section with a discussion on content editing and how a browser extension can modify text or social metrics. Section 5 offers a detailed description of how these manipulations can be accomplished in practice.
@@ 3.1. Down-ranking
The ability to down-rank content allows researchers to investigate alternative moderation strategies. Harmful content can be dynamically penalized in the ranking, making it available only when the user is interested in consuming more content and scrolling further down their feed. Given criteria for down-ranking content, such interventions are generally easy to implement. The extension should select the posts that must be penalized and estimate how far down the feed the item ought to be down-ranked. The new position can be determined based on a fixed offset or based on the score of the content. For example, researchers may decide to add 100 positions to all down-ranked posts or use a content-based offset, making sure that very harmful posts receive less exposure. In practical terms, the extension should keep track of the number of posts consumed by the user and insert the post when the browser’s viewport reaches the desired position. This step can be achieved easily on most platforms by inserting the same content removed in the previous position. This intervention offers valuable insights into the impact of ranking. It can provide concrete recommendations for the social media platform on adapting the algorithm to penalize problematic content and achieve the desired outcome. This intervention allows testing interventions that can reduce exposure to some content but do not entirely remove the recommended posts–a topic that may be sensitive for those opposing moderation strategies.
As an alternative, an extreme form of down-ranking is the complete removal of specific content. This logic, similar to how ad blockers work, operates on the definition of an exclusion logic or metric that guides the extension in identifying the posts that should be removed from the user’s feed. This intervention can provide valuable insights into the impact of filtering out posts with particular characteristics. This approach allows researchers to investigate research questions such as the impact of stricter moderation methods by reducing the content that has the potential to polarize communities, affect people emotionally, or to investigate the impact of alternative feed designs such as removing all the posts with videos or images.
@@ 3.2. Up-ranking
Without access to the complete platform inventory, up-ranking content in the feed presents more challenges than down-ranking content. The content returned by the server is obtained by scoring the inventory, which includes the entire candidate set of potential posts available for display. Without the full set or a meaningful approximation, the intervention can operate only on the pre-selected content delivered to the user, representing the top items the curation algorithm has already prefiltered. For example, when the platform’s algorithm selects content based on engagement, the posts available for re-ranking may consist only of high-engagement posts. Access to the full inventory could significantly expand the dynamic range of content available for re-ranking, potentially showing a more prominent effect of an intervention. However, compiling the full or even partial inventory may require more active support from the platforms or data access agreements, which may be difficult or expensive to set up. Possible approaches include up-ranking posts loaded by the website but not visible in the browser’s viewport, pre-fetching content with the extension by simulating scrolling or inserting entirely new posts from the platform inventory obtained in other ways.
@@
Inserting new posts provides a mechanism for customizing the feed and exposing the user to content not selected by the curation algorithm. This intervention can help researchers investigate research questions that rely on adding content blended in the context of the original feed. Some examples include adding more content promoting positive emotions, adding out-party posts to counter filter bubbles, or adding posts on a specific topic to change the topic distribution of the feed. This strategy is particularly helpful when the type of posts the researchers are interested in investigating is rare, and the goal is to ensure the feed has enough relevant content.
@@
From a technical point of view, this approach requires careful consideration. The inserted posts could be created in three ways: (i) generating entirely new posts that fit the requirements of the experiment, (ii) monitoring posts by a curated list of accounts that are likely to post suitable content, and (iii) transferring a post that matches these requirements from another feed. The first approach may present challenges because it requires replicating a fully functional social media post, including user interactions, such as likes, comments, and shares. This method may allow full customization but demands careful attention to detail to ensure the post supports real-world user interaction. Failing to recreate all these interactions accurately (i.e., the share button not working) may disrupt the user experience and compromise the study’s validity by influencing user behavior in unanticipated ways.
@@
The second approach is monitoring a set of public accounts that are likely to post content that fits the requirements of the intervention. This method requires running a background process that continuously collects and scores the content posted by the selected accounts. There are two challenges with this approach: (i) the content that individual users post varies and curating a list of suitable accounts effectively may be challenging for some interventions; and (ii) the data collection may require additional data access, which may be prohibitively expensive.
@@
The third approach is transferring content from other existing feeds. This approach offers a more pragmatic alternative, allowing researchers to utilize authentic posts. This technique has a downside: researchers must implement a mechanism to find or produce posts that qualify for the intended intervention, especially without full inventory. This can be done by drawing the post from a dedicated account, using tweets from the same user previously removed, or monitoring the feeds of other participants in the experiment. In the former case, this method requires researchers to ensure that the content selected for transfer respects user privacy and does not inadvertently expose private or sensitive information. Finally, researchers employing this intervention must reflect on the potential impact of context when transferring posts between feeds, as the relevance and reception of a post may vary significantly based on its original context.

### v2

@@ Abstract.
Social media plays a central role in shaping public opinion and behavior, yet performing experiments on these platforms and, in particular, on feed algorithms is becoming increasingly challenging. This guide offers practical recommendations for researchers developing and deploying field experiments focused on real-time reranking of social media feeds. The article is organized around two contributions. First, we provide an overview of an experimental method using web browser extensions that intercepts and reranks content in real time, enabling naturalistic reranking field experiments. We then describe feed interventions and measurements that this paradigm enables on participants’ actual feeds, without requiring the involvement of social media platforms. Second, we offer concrete technical recommendations for intercepting and reranking social media feeds with minimal user-facing delay, and provide an open-source implementation. This document aims to summarize lessons learned in running field experiments on social media, provide concrete implementation details, and foster the ecosystem of independent social media research. Finally, we release the source code that serves as a blueprint for implementing future feed-ranking experiments: https://github.com/StanfordHCI/FeedMonitor.
@@ 1. Introduction
Figure 1. Diagram of the typical feed reranking steps using a browser extension. First, the extension identifies the posts that need to be rescored (optional if all posts are included). The ranking model assigns a new score to the posts, and the updated data is finally displayed in the browser. The modified feed can be enriched with custom widgets, such as surveys, to assess the intervention’s impact.
@@
Social media plays a critical role in forming public opinion and influencing both individual behaviors and societal outcomes (Weeks et al., 2017). Many rely on social media to follow the news and stay informed (Center, 2023). Central to this media consumption behavior is the feed, which selects and ranks the social media posts users see. It is the primary interface through which platforms direct users’ attention and users engage with content. Most major platforms, including Facebook, X (Twitter), and Reddit, rely on their feeds to provide users with a centralized view of relevant content. However, in doing so, these platforms must make algorithmic curation choices (Backstrom, 2016). In response, researchers have called for special attention to the impact of these feed algorithms, for example, the kinds of content and behaviors that they up-rank or down-rank (Allen and Tucker, 2025). These design choices are of substantial interest because they may profoundly shape our experiences and beliefs. Of particular concern are feed algorithms that aim to maximize user engagement and inadvertently amplify undesirable content, which may exacerbate polarization (Rathje et al., 2021; Kubin and von Sikorski, 2021), misinformation (Valenzuela et al., 2019), and antisocial behaviors (Saurwein and Spencer-Smith, 2021).
@@
Investigating the causal impact of the feed ranking in an ecologically valid way is challenging. Despite efforts to open-source some of the key components of the platforms’ ranking algorithms (Twitter, 2023), their “black box” nature, reinforced by the lack of data (Tromble, 2021), complicates attempts to evaluate their causal effects on individual users and society as a whole. While the “logic” behind the algorithm may be open source, the algorithms’ outputs inherently depend on closed machine learning models that score aspects of each post and determine what is visible in the feed and in what order. Another challenge is the increasingly limited access to the users’ post inventory, i.e., the posts that would be considered to be shown in the users’ feeds. Incomplete access to the key components of the feed—algorithms, models, and inventory—inherently limits the validity of any investigation.
@@
This article aims to bridge this gap by offering a practical guide for conducting independent field experiments to investigate the impact of feed ranking in social media, building on our own experience in designing and deploying such interventions (Piccardi et al., 2025). The guidelines focus on facilitating the design of experiments that do not require direct cooperation with the social media platform, enabling audits of feed algorithms and the development of interventions to improve the design of online social spaces. We describe how to implement experiments on the web interfaces of social media platforms using browser extensions, a feature available in all major browsers, by intercepting and customizing the network requests that populate the feed. We focus on specific interventions that this method supports, including feed reranking and content editing, and describe the relative advantages and disadvantages of each. We do not aim to exhaustively list all possible ways to implement a field experiment, but we do provide concrete building blocks for designing custom experiments.
@@
Figure 1 summarizes the main steps of a typical field experiment designed to customize the ranking of a social media feed with a browser extension. This guide focuses on designing interventions that do not affect the users’ typical experience on the platform, aiming to limit latency and any interface behavior that can disrupt the user experience. To achieve this goal, the browser extension (1) captures the feed, (2) applies any experimental transformations, and then (3) inserts any custom user interface elements. First, it reads the feed to identify posts that need editing based on the experiment’s objective. For example, the experiment may focus on all the content in the feed, or it may target specific posts, such as politics or misinformation. Then, the extension can apply the desired intervention to the selected posts (or newly inserted posts), such as changing their ranking, rewriting their text, or editing their social metrics. Finally, the extension can change the webpage’s design or add widgets that enhance the interface and enable new user interactions. Such modifications enable researchers to design custom user interactions or assess the effect of a specific intervention by adding targeted surveys.
@@
This article is organized as follows. In Section 2, we summarize the state-of-the-art and highlight studies closely related to our work. In Section 3, we present opportunities and challenges in changing the feed ranking with a browser extension, followed by an overview of possible approaches to measuring the outcomes of the field experiment. Finally, Section 5 provides key implementation details to develop the feed experiment, ranging from ensuring low-effort user registration to intercepting and editing the feed. We conclude with a discussion of implications. With this article, we release an open-source implementation of the proposed framework (StanfordHCI, 2026).
@@ 2. Background
In this section, we summarize previous efforts to study social media through experiments and position this article in the current literature. On-platform feed experiments are the gold standard for inferring the causal effects of interventions on social media, as they test interventions in the same setting where users would typically encounter them, maximizing ecological validity. An alternative to running experiments on the platform is to create experimental clones of the platform and simulate a social media environment (Voggenreiter et al., 2023; Epstein and Lin, 2022; Jia et al., 2024). Developing a custom social media feed simulator is a meaningful approach, as it allows greater control over the experimental setup, including the posts the user sees and the features offered by the websites. This full control over the feed has significant advantages, as participants can be exposed to the same content, potentially leading to a more controlled experiment that can help measure small effects. However, the downside is that participants’ online experiences are generally hard to reproduce, and the generalizability of the findings may be limited. Participants may be exposed to a different interface and a different type of posts they typically consume, miss realistic social signals and dynamics, or be required to engage in activities not aligned with their daily habits. The requirements to engage with a custom platform may also exacerbate phenomena such as the Hawthorne effect, potentially altering participants’ behavior in ways that do not fully generalize to a real-world scenario.
Previous work shows the effectiveness of field experiments on social media and that they can bring valuable findings. Some examples include investigating the effect of changing Facebook’s feed ranking from engagement-based to chronological during a US election (Guess et al., 2023), the impact and propagation of emotions (Kramer et al., 2014), the amplification of political discourse (Huszár et al., 2022), the effect of perspective-taking on political polarization (Saveski et al., 2022), or using ads to reduce misinformation (Lin et al., 2024) and support charitable donations (Jilke et al., 2019).
However, designing a field experiment with real-time intervention on social media feeds poses many technical challenges, especially when conducted without the platforms’ involvement (Allen and Tucker, 2025; Mosleh et al., 2022; Guess, 2021). The present work aims to support researchers with practical recommendations for designing field experiments focused on feed ranking.
@@
While many field experiments of interest only require changing participants’ feeds, some experiments remain out of scope for this approach. For example, this approach cannot be used to perform network-based randomized experiments or other experimental protocols that require changing entire groups of users’ experiences together (Ugander et al., 2013; Saveski et al., 2017), e.g., to understand equilibrium effects rather than individual effects. Our approach also cannot easily access the user’s full inventory unless the experimenter implements a separate application layer to acquire it directly from the platform. In this paper, we will focus on changes to the feed items that have already been selected to be shown to the user.
Concrete Use-case. The effectiveness of the described framework was demonstrated in a large-scale experiment on X involving 1,256 participants (Piccardi et al., 2025). Over a 10-day period leading up to the 2024 US presidential election, participants’ feeds were reranked to evaluate the impact of content exposure on affective polarization and emotional responses. Participants’ posts were processed by a backend system, initially filtered for political content using a custom BERT classifier. Political posts were then scored by GPT using well-established social science constructs to estimate each post’s anti-democratic content. Posts were subsequently up-ranked or down-ranked according to the assigned experimental condition. Compared to a control group whose feeds remained unaltered, participants exposed to fewer posts that contained anti-democratic content reported decreased affective polarization and a temporary reduction in anger and sadness. Conversely, increased exposure to anti-democratic content posts resulted in heightened affective polarization and negative emotions. Given the urgency of independent research on the impact of social media platforms (Allen and Tucker, 2025), here we summarize lessons learned from implementing this experiment to support the community in designing similar studies.
@@ 2.1. Key concepts
Browser extensions. This guide focuses on modifying social media feeds using Browser extensions. Browser or web extensions are add-ons typically available via the browser-specific store, designed to enhance the browser’s functionality. They support developers in customizing the browser’s behavior and modifying websites’ content. This capability is advantageous for developing custom social media feed interventions, enabling design-control trials or A/B tests. From the perspective of the experimental participants, the only technical prerequisite is to install the extension and accept the permissions. With authorization, browser extensions can modify the pages of interest to achieve the desired intervention. These modifications to the web pages range from updating the content and its visual appearance to the website’s behavior, such as the response to user interactions. Section 5 provides an overview of implementation details.
Platform Inventory. In this guide, we refer to the full database of candidate posts used by the feed algorithm as the participant’s inventory. This is a crucial concept because, depending on the platform of interest, the level of access to the candidate set can enable or limit some interventions. When the feed is organized chronologically, the algorithm can sort the posts of relevant users (e.g., friends or followed users) by time and return the most recent ones. When the feed is algorithmically curated, the algorithm must evaluate a large set of posts from users within and outside the focal user’s first-degree social network. This process typically relies on multiple steps. For instance, Twitter/X’s “For you” algorithm, as described in the open-sourced implementation (Twitter, 2023), first selects content using a high-recall retrieval system that scans a large amount of data, which is then scored and ranked by a more refined model.
@@ 3. Feed Ranking
This section summarizes the potential of browser extensions to customize social media feeds. The reranking can be a complete shuffle of the feed based on a new objective or more focused, for example, up-ranking or down-ranking content that the researchers or users (Kolluri et al., 2026) believe should receive more or less exposure. In this guide, we focus on these two use cases, up-ranking and down-ranking, where the content of interest is made easier or harder to reach, respectively, but these recommendations can be adapted for different types of rank adaptations. These edits can be achieved by combining up-ranking and down-ranking operations, which constitute the building blocks of a feed intervention.
A key functionality of the system handling the experiment is identifying relevant content and scoring it according to the metric of interest. Depending on the research goal, this capability can be achieved at different levels of complexity, ranging from simple keyword matching to more complex logic that relies on external AI services, such as large language models (LLMs). For example, feed posts can be scored and rearranged to make hostile political content harder to reach, or to make posts that reflect positive emotions easier to reach.
These manipulations offer a unique opportunity to investigate the causal effects of variations of the rank objective in a realistic scenario. We summarise the challenges and limitations of both implementations. Additionally, a browser extension can edit post content. We conclude this section with a discussion on content editing and how a browser extension can modify text or social metrics. Section 5 offers a detailed description of how these manipulations can be accomplished in practice.
@@ 3.1. Down-ranking
The ability to down-rank content allows researchers to investigate alternative moderation strategies. Potentially harmful content can be dynamically penalized in the ranking, making it available only when the user is interested in consuming more content and scrolling further down their feed. Given the criteria for down-ranking content, such interventions are generally easy to implement. The extension should select the posts that must be penalized and estimate how far down the feed each item should be down-ranked. The new position can be determined based on a fixed offset or based on the score of the content. For example, researchers may decide to push all down-ranked posts 100 positions lower in the feed or use a content-based offset to ensure that very harmful posts receive less exposure. In practice, the extension should keep track of the number of posts the user has consumed and insert a post when the browser’s viewport reaches the desired position. This step can be easily achieved on most platforms by inserting the same content that was removed in the previous position. This intervention offers valuable insights into the impact of ranking. It can provide concrete recommendations for the social media platform on how to adapt its algorithm to penalize problematic content and achieve the desired outcome. This intervention allows testing interventions that can reduce exposure to some content without entirely removing the recommended posts—a topic that may be sensitive for those opposing moderation strategies.
As an alternative, an extreme form of down-ranking is the complete removal of specific content. This logic, similar to how ad blockers work, operates on the definition of an exclusion logic or metric that guides the extension in identifying the posts that should be removed from the user’s feed. This intervention can provide valuable insights into the impact of filtering out posts with particular characteristics. This approach allows researchers to investigate questions such as the impact of stricter moderation methods that reduce content with the potential to polarize communities or emotionally affect people, and the impact of alternative feed designs, such as removing all posts with videos or images.
@@ 3.2. Up-ranking
Without access to the full platform inventory, up-ranking content in the feed poses more challenges than down-ranking it. The content returned by the server is obtained by scoring the inventory, which includes the entire candidate set of potential posts available for display. Without the full set or a meaningful approximation, the intervention can operate only on the pre-selected content delivered to the user, representing the top items the curation algorithm has already prefiltered. For example, when the platform’s algorithm selects content based on engagement, the posts available for reranking may consist only of high-engagement posts. Access to the full inventory could significantly expand the dynamic range of content available for reranking, potentially showing a more prominent effect of an intervention. However, compiling the full or even partial inventory may require more active support from the platforms or data access agreements, which may be difficult or expensive to set up. Possible approaches include up-ranking posts loaded by the website but not visible in the browser’s viewport, pre-fetching content with the extension by simulating scrolling, or inserting entirely new posts from the platform inventory obtained in other ways.
@@
Inserting new posts provides a way to customize the feed and expose users to content not selected by the curation algorithm. This intervention can help researchers investigate research questions that rely on adding content blended into the original feed. Some examples include adding more content that promotes positive emotions, adding out-party posts to counter filter bubbles, or adding posts on a specific topic to shift the feed’s topic distribution. This strategy is particularly helpful when the type of posts the researchers are interested in investigating is rare, and the goal is to ensure the feed has enough relevant content.
@@
From a technical standpoint, this approach requires careful consideration. The inserted posts can be sourced in three ways: (i) generating entirely new posts that fit the requirements of the experiment, (ii) monitoring posts by a curated list of accounts that are likely to post suitable content, and (iii) transferring a post that matches these requirements from another feed. The first approach may present challenges because it requires replicating a fully functional social media post, including user interactions, such as likes, comments, and shares. This method may allow full customization but demands careful attention to detail to ensure the post supports real-world user interaction. Failing to recreate all these interactions accurately (i.e., the share button not working) may disrupt the user experience and compromise the study’s validity by influencing user behavior in unanticipated ways.
@@
The second approach is to monitor a set of public accounts likely to post content that meets the intervention’s requirements. This method requires running a background process that continuously collects and scores the content posted by the selected accounts. There are two challenges with this approach: (i) the content that individual users post varies, and curating a list of suitable accounts effectively may be challenging for some interventions; and (ii) the data collection may require additional data access, which may be prohibitively expensive.
@@
The third approach is to transfer content from existing feeds. This approach offers a more pragmatic alternative, allowing researchers to utilize authentic posts. This technique has a downside: researchers must implement a mechanism to find or produce posts that qualify for the intended intervention, especially without a full inventory. This can be done by drawing the post from a dedicated account, using posts from the same user previously removed, or monitoring the feeds of other participants in the experiment. In the former case, this method requires researchers to ensure that the content selected for transfer respects user privacy and does not inadvertently expose private or sensitive information. Finally, researchers employing this intervention must reflect on the potential impact of context when transferring posts between feeds, as the relevance and reception of a post may vary significantly based on its original context.